# Massimiliano Ciranni1 These authors contributed equally to this work.    Vito Paolo Pastore1,2∗    Roberto Di Via1∗    Enzo Tartaglione3    Francesca Odone1    Vittorio Murino2,4

### v2

@@ 3.2 Diffusing the Bias
In the context of mitigating bias in classification models, the tendency of a CDPM to approximate the per-class biased distribution represents a key feature for training a bias amplification auxiliary model.
Figure 2: Schematic representation of our DDB framework. The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract bias signals from real images. These signals guide debiasing recipes (i.e., two-step and end-to-end methods) that process the original training data to produce a debiased model.
Diffusion Process Background.
The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [31]. Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$. This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$. The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\boldsymbol{\epsilon}_{\theta}$:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (1)
@@
In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes. Classifier-Free Guidance (CFG) [32] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
The CFG technique randomly omits conditioning during training (e.g., with probability $p_{\text{uncond}}=0.1$), enabling the model to learn both generation modalities. During the sampling process, a guidance scale $w$ modulates the influence of conditioning. When $w=0$, the model relies solely on the conditional model. As $w$ increases ($w\geq 1$), the conditioning’s effect is intensified, potentially resulting in more distinct features linked to $y$, thereby increasing fidelity to the class but possibly reducing diversity, whereas lower values help to preserve diversity by decreasing the influence of conditioning. The guided noise prediction is given by:
$\boldsymbol{\epsilon}_{t}=(1+w)\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)-w\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (2)
where $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)$ is the noise prediction conditioned on class label $y$, and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction. This modified noise prediction replaces the standard $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ term in the reverse process formula (Equation 1). In this work, we empirically show how the CDPM learns and amplifies underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation. In the next Section, we describe how such synthetic images can be used to train a robust Bias Amplification model.
@@ 3.3 DDB: Bias Amplifier and Model Debiasing
As stated in Sec. 2, a typical unsupervised approach to model debiasing relies on an intentionally biased auxiliary model, named here as Bias Amplifier (BA). This model can be exploited in either two-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
3.3.1 Recipe I: two-step debiasing
Figure 3: Overview of Recipe I’s two-step debiasing approach.
As a two-step, we adopt the bias supervised algorithm G-DRO [8] as a proven technique for the pure debiasing step. Instead of using the real bias labels (considered unknown), we estimate pseudo-labels performing an inference step by feeding the trained BA with original training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified. Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on top of the BA misclassifications. Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate it as
@@ 3.3.1 Recipe I: two-step debiasing
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA model on a real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Equation 3). Once training subgroups pseudo-labels are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3. This filtering operation refines the plain error set, restricting bias-conflicting selection to the hardest training samples, with potential benefits for the hardest correlation settings. In our experiments, we provide an ablation study comparing filtering and plain error set alternatives, in terms of performance.
@@ 3.3.2 Recipe II: end-to-end debiasing
A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] intentionally biased auxiliary models. Here, we design an end-to-end debiasing recipe, denoted as Recipe II, incorporating our Bias Amplifier by customizing a widespread general scheme, introduced in Learning from Failure (LfF) [6]. This approach leverages an intentionally biased model trained using the GCE to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score. Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4. Such loss function is used to obtain a weighting factor for the target model’s loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$. Intuitively, $r$ is low for bias-aligned and high for bias-conflicting samples.
@@ 4 Experiments
In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis. In the Suppl. Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
@@ 4.1 Datasets
Our experiments utilize four distinct datasets, a synthetic one, Corrupted CIFAR-10 (CIFAR10-C), and three real-world datasets, namely Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), and Biased Action Recognition (BAR). Specifically, CIFAR10-C [33] is a synthetic dataset derived from CIFAR-10, altered with various corruptions (such as noise, blur, and pixelation): we utilize the configurations with  $\rho=\{0.950,0.980,0.995\}$, the same used in [34]. Waterbirds [35] is an image dataset with inherent biases, exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs. land). BFFHQ [36] builds upon FFHQ [37] by introducing demographic biases for face recognition  ($y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$). Finally, BAR [6] dataset is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts.
@@ 4.2 Implementation Details
In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels. For computational efficiency, input images are resized to $64\times 64$ pixels (except for Corrupted CIFAR-10 whose original size is $32\times 32$, and thus kept as is). Input images are normalized with respect to ImageNet’s mean and standard deviation. The CDPM is trained from scratch for $70,000$, $100,000$, $150,000$, and $200,000$ iterations for the datasets BAR, CIFAR10-C ($\rho=0.950$), Waterbirds, and BFFHQ, respectively. Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$). Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR [38] scheduler with a warm-up period for the first $10\%$ of the total iterations. As per the Bias Amplifier training, we use synthetic images from CDPM. A Densenet-121 [39] with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head. The training uses regular Cross-Entropy loss function and the AdamW optimizer [40]. For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds and Corrupted CIFAR-10. Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs. We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips. For our debiasing recipes we rely on the same backbones used in the existing literature, i.e.ResNet-18 for Corrupted CIFAR-10, BFFHQ, and BAR, and a ResNet-50 for Waterbirds [6, 4, 25]. In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments. Finally, similarly to [20, 6, 3], we do not rely on a validation set for regularization. This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization [10].
@@ 4.3.1 Synthetic bias-aligned image generation
First, we present qualitative image generation results across multiple datasets (see Fig. 5), to validate the effectiveness of our bias diffusion approach. The generated samples maintain good fidelity and capture the typical bias patterns of each training set, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases. This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample. In addition, we also carried out an objective evaluation of the generated image content by adopting an image captioner, specifically BLIP-2 [41]. We estimate a caption for each image per class and count the frequency of every word. Suppl. Material (Sec. A.Bias Validation via Captions in Synthetic Images) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced. The same behavior also emerges in other dataset used in this study (reported in Sec. A.Bias Validation via Captions in Synthetic Images), providing a good description of the bias attributes. These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
4.3.2 Debiasing accuracy
Method Corrupted CIFAR-10
$\rho=0.995$ $\rho=0.980$ $\rho=0.950$
[…]
DDB-I (ours) $\bf 31.46\scriptsize{\pm 1.08}$ $\bf 44.90\scriptsize{\pm 1.17}$ $53.38\scriptsize{\pm 0.51}$
[…]
Table 1: Results of DDB Recipes I and II on Corrupted CIFAR-10 and available unsupervised debiasing state-of-the-art works. The best results are highlighted in bold, second-best are underlined.
Tables 1 and 2 report the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5]. For BFFHQ, we report the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20]. BAR does not provide bias-attribute annotations, hence we report just the average test accuracy, in line with former works (e.g., [6]). For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. For Corrupted CIFAR-10 (Table 1), DDB-I and DDB-II provide the best performances overall, especially on the most challenging $\rho$ settings, providing a maximum improvement of $1.51\%$ for $\rho=0.995$, $3.12\%$ for $\rho=0.980$, and $4.67\%$ for $\rho=0.950$, all with respect to the DFA method [34]. Notably, both recipes outperform the state-of-the-art performances, showing a stable behavior.
Method Unsup Waterbirds BAR BFFHQ
WGA Avg. Confl.

### v3

@@ 3.2 Diffusing the Bias
In the context of mitigating bias in classification models, the tendency of a CDPM to approximate the per-class biased distribution represents a key feature for training an auxiliary bias amplified model.
The Diffusion Process.
The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [32]. Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$. This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$. The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\bm{\epsilon}_{\theta}$:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (1)
@@
In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes and semantics. Classifier-Free Guidance (CFG) [33] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
The CFG technique randomly omits conditioning during training (e.g., with probability $p_{\text{uncond}}=0.1$), enabling the model to learn both generation modalities. During the sampling process, a guidance scale $w$ modulates the influence of conditioning. When $w=0$, the model relies solely on the conditional model. As $w$ increases ($w\geq 1$), the conditioning effect is intensified, potentially resulting in more distinct features linked to $y$, thereby increasing fidelity to the class while possibly reducing diversity, whereas lower values help to preserve diversity by decreasing the influence of conditioning. The guided noise prediction is given by:
$\bm{\epsilon}_{t}=(1+w)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)-w\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (2)
where $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)$ is the noise prediction conditioned on class label $y$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction. This modified noise prediction replaces the standard $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ term in the reverse process formula (Equation 1). In this work, we empirically show how CDPM learns and amplifies the underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation.
@@ 3.3 DDB: Bias Amplifier and Model Debiasing
As stated in Sec. 2, a typical unsupervised approach to model debiasing relies on an auxiliary intentionally-biased model, named here as Bias Amplifier (BA). This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
3.3.1 Recipe I: 2-step debiasing
Figure 3: Overview of Recipe I’s 2-step debiasing approach.
The adopted 2-step approach consists in 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) apply a bias supervised method to train a debiased target model for classification. For the latter, we use the group DRO algorithm [8] (G-DRO) as a proven technique for the pure debiasing step. In other words, being in the unsupervised bias scenario where the real bias labels are unknown, we estimate bias pseudo-labels performing an inference step by feeding the trained BA with the original actual training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified. Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on the BA misclassifications. Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate bias-conflicting samples as
@@ 3.3.1 Recipe I: 2-step debiasing
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq. 3). Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3.
The above filtering operation refines the plain error set, restricting bias-conflicting sample selection to the hardest training samples, with potential benefits for the most difficult correlation settings ($\rho>0.99$). Later in the experimental section, we provide an ablation study comparing different filtering ($\gamma$) configurations and plain error set alternatives.
@@ 3.3.2 Recipe II: end-to-end debiasing
A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] auxiliary intentionally-biased models. Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method [6]. LfF leverages an intentionally-biased model trained using Generalized CE (GCE) loss to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score. Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4. Such loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$. Coarsely speaking, $r$ should be low for bias-aligned and high for bias-conflicting samples.
@@ 4 Experiments
In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis.
In the Supplementary Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
@@ 4.1 Datasets
Our experiments utilize five distinct datasets: Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), Biased Action Recognition (BAR), and ImageNet-9/ImageNet-A. Waterbirds  [8] is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs. land). It has 4,795 training images, 1,199 images for validation, and 5,794 for testing. BFFHQ [4] builds upon FFHQ [34] by introducing demographic biases for face recognition, while BAR  [6] is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts. For these two last datasets, we specifically refer to their original versions introduced respectively in [6] and [4]: BFFHQ has a total of 19,200 training images where  ( $y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$). Then, it provides 1,000 validation images (all bias-aligned) and 1,000 testing images with uniformly distributed labels and bias attributes. BAR presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing. Finally, ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions. Originally introduced by Hendrycks et al. [35] to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 [36], a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases [36]. As such, both our CDPM and Bias Amplifier are trained on ImageNet-9. Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of [36, 1, 31]
@@ 4.2 Implementation Details
In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels. For computational efficiency, input images are resized to $64\times 64$ pixels. Input images are normalized with respect to ImageNet’s mean and standard deviation. The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, and ImageNet-9 respectively. Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$). Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR [37] scheduler with a warm-up period for the first $10\%$ of the total iterations. As per the Bias Amplifier training, we use synthetic images from CDPM. A Densenet-121 [38] with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head. The training uses regular Cross-Entropy loss function and the AdamW optimizer [39]. For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9. Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs. We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips. For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds [6, 4, 23]. In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments. Finally, similarly to [20, 6, 3], we do not rely on a validation set for regularization. This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization [10].
@@ 4.3.1 Synthetic bias-aligned image generation
First, we present qualitative image generation results across multiple datasets (see Fig. 5), to validate the effectiveness of our bias diffusion approach. The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases. This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample. In addition, we also carried out an objective evaluation of the generated image content by adopting an image captioner, specifically BLIP-2 [40]. We estimate a caption for each image per class and count the frequency of every word. Suppl. Material (Sec. A.E) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced. The same behavior also emerges in other datasets used in this study, where a consistent description of the bias attributes can be noted, as reported in Sec. A.E. These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
4.3.2 Classification accuracy
Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5]. For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20]. BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy. For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. It’s worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
Method Unsup Waterbirds BAR BFFHQ ImageNet-A
WGA Avg. Avg. Confl. Avg.